OpenStudy (calculusxy):

MEDAL! The radius of circle B is 2.5 times the radius of circle A. If the area of circle A is 8 square yards, what is the area of circle B?

OpenStudy (studygurl14):

First step. Find the radius of circle A using the area of 8 yd^2

OpenStudy (calculusxy):

Well here is the work that I did.

OpenStudy (studygurl14):

Okay.

OpenStudy (calculusxy):

\[3.14 \times r^2 = 8\] \[r^2 = 3\] \[\sqrt{3}=1.73\]

OpenStudy (calculusxy):

\[1.73 \times 2.5 = 4.33 yards\]

OpenStudy (studygurl14):

First off, I wouldn't round 2.548 to 3...

OpenStudy (calculusxy):

Okay

OpenStudy (studygurl14):

it's best to keep it as accurate as possible, and only round towards the end.

OpenStudy (calculusxy):

understood..

OpenStudy (studygurl14):

So take the square root of 2.5477. What do you get?

OpenStudy (studygurl14):

round to the nearest thousandth

OpenStudy (calculusxy):

1.568

OpenStudy (calculusxy):

And then i multiply by 3 right?

OpenStudy (studygurl14):

i got something different...

OpenStudy (calculusxy):

Sorry 2.5

OpenStudy (studygurl14):

no...i got 1.596, not 1.568

OpenStudy (calculusxy):

yes. i typed it wrong in the calculator.

OpenStudy (calculusxy):

and then what do i do?

OpenStudy (studygurl14):

That's the approx. radius circle A. Multiply by 2.5 to find the radius of circle B

OpenStudy (calculusxy):

@StudyGurl14 Thank you.